# Detectability of dark matter subhalo impacts in Milky Way stellar streams

### v2

@@ 5.1 Stream ranking
Figure 18: Visualized stream catalog ranked by detectability under Via + Gaia + LSST10 (left), and their sky plots (right).
Figure 19: Visualized stream catalog, where color scale indicates the the minimum detectable subhalo mass, $M_{\mathrm{sh}}^{\mathrm{min}}$, with respect to the best observational scenario. The stream properties of stellar density $\lambda$, width $\sigma_{\theta}$, and distance to observer $r_{0}$ all play a role in determining $M_{\mathrm{sh}}^{\mathrm{min}}$. The five streams with lowest $M_{\mathrm{sh}}^{\mathrm{min}}$ are labeled.
Figure 20: Improvement from DESI + Gaia to Via + Gaia + LSST in $\log_{10}M_{\mathrm{sh}}^{\mathrm{min}}/\mathrm{M}_{\odot}$ as functions of stream width (left), stream distance (middle) and stream density (right). The color scale indicates the the minimum detectable subhalo mass, $M_{\mathrm{sh}}^{\mathrm{min}}$, with respect to the best observational scenario.
@@ 6.1 Generalization to realistic streams
Circular orbits and observer location – We expect that some of the largest differences between our analysis and more realistic models will arise in cases of large variation of heliocentric and/or Galactocentric distance along the stream orbit. In our study, streams are moving on circular orbits and observers are located at the Galactic Center. Accounting for eccentric orbits and placing the observer at the Sun will change the shape of subhalo-induced perturbations, their time dependence, as well as the relative importance of different observables. For a stream with a large variation in heliocentric distance along its length, a subhalo impact will be compressed to a shorter angular arc from the point of view of the observer, leading to a diminished detectability. Subhalo impacts will also be compressed and stretched depending on whether impacts occurred closer to pericenter or apocenter, as well as whether the stream today is observed near pericenter or apocenter. This will lead to much stronger dependence of subhalo detectability on impact time. Furthermore, there will be a stronger requirement on how well we can model the orbit of the progenitor, since observables are defined with respect to a well-understood smooth stream track. We expect streams with relatively constant heliocentric and Galactocentric distances along the stream to have the most similar detectability as our idealized study.
@@
Milky Way potential – The actual Milky Way potential is much more complicated than a simple logarithmic potential used here. Baryonic structures like the bar, spiral arms, as well as giant molecular clouds (GMC) and globular clusters (GC), could potentially cause perturbations in the streams. This will further raise the threshold of detectability, beyond the ideal case shown here. At the same time, kinematic information like radial velocity and proper motions could help distinguish between the impact of baryonic structures and DM subhalos (Bonaca et al., 2020; Hilmi et al., 2024; Price-Whelan et al., 2016).
@@
Stream model – Our stream model uses position/velocity dispersions assuming a particle spray model, where the parameters governing the dispersions are taken from Fardal et al. (2015). We investigated another particle spray model, based on a newer suite of simulations in Chen et al. (2024). This model includes initial radial velocity dispersions for the released particles and generates a stream with twice the $\sigma_{v_{r}}$ and $\sigma_{v_{t}}$ compared to a stream generated using Fardal et al. (2015), holding the angular width $\sigma_{\theta}$ fixed. We found that this difference does not change the estimated detectability in a significant way, because the shift in the $z$ direction $\overline{\Delta\theta}$ is the most effective observable in subhalo detection (see Fig. 9). The dispersion for this observable is governed primarily by the directly observed quantity $\sigma_{\theta}$. In addition to affecting the dispersions, the dynamics of progenitor disruption could induce perturbations in the streams and further raise the threshold for detectability. However, these effects are expected to mainly be in the region near the progenitor.
@@
Rather than fixing $r_{s}(M_{\mathrm{sh}})$, the degeneracy in mass and velocity can also be broken by using other subhalo potentials (such as Hernquist or NFW) and performing the full orbit integration in presence of the DM subhalo. To assess the impact of this choice, we used an orbit integration method where we instead model the perturbed stream using the perturbed orbit of the progenitor. The details are discussed in App. B. We then evaluated how expected confidence intervals change when the $r_{s}(M_{\mathrm{sh}})$ relationship is removed and/or the Hernquist potential is used. For the Hernquist potential, the CDM mass-radius relation follows (Erkal et al., 2016):
@@
Fig. 21 shows a comparison using orbit integration with our default approach (Analytic + Plummer + $r_{s}(M_{\mathrm{sh}})$, see also Fig. 21). In all cases, we evaluated the test statistic on the Asimov data set. We find that using orbit integration generally results in tighter confidence intervals, even when $r_{s}$ is a free parameter of the fit. This indicates that our assumption of $r_{s}(M_{\mathrm{sh}})$ is not overly restrictive, and in fact the model likely underestimates the detectability compared to orbit integration. The results also show that Plummer and Hernquist potentials give similar signal strengths, as long as consistent $r_{s}$ values are used (see also the lower right panel of Fig. 26). Assessing the impact on the minimum detectable subhalo mass is much more expensive computationally, since it requires many more statistical realizations, and therefore we reserve this type of analysis for dedicated studies of individual streams.
Figure 21: Comparison of the test statistic $-t$, as defined in Eq. 28, using the analytic model and using an orbit integration method. The dataset is generated from $M_{\mathrm{sh}}^{\mathrm{true}}=10^{8}~{}\mathrm{M}_{\odot}$. The stream has properties similar to GD-1 ($\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~{}\mathrm{kpc}$, $\lambda=100~{}\mathrm{deg}^{-1}$) and the observational scenario is LSST 10 year sensitivity.
More generally, the orbit integration method can be used to account for many of the concerns above. This would allow for a more general Milky Way potential, including known structures like the bar, spiral arms, and the Large Magellanic Cloud. It also provides an intermediate approach to modeling perturbed streams, which avoids orbit integration of the entire stream. It is a good approximation for CDM subhalos and again Gaussian noise can be added to account for dispersion. However, given that this approach is still much slower than the analytic model, we reserve this for future work.
@@ 6.2 Comparisons with previous work
Our work draws on Erkal & Belokurov (2015b), which derived the analytic model. That work focused on parameter estimation from a subhalo impact rather than the minimum detectable subhalo. Using a Bayesian framework, the 6 observables from the analytic model are used to infer all the parameters of the subhalo impact at the same time. To address the subhalo velocity-mass degeneracy of the analytic model discussed above, they imposed a subhalo velocity prior rather than a mass-radius relationship. The confidence bands are then highly dependent on the velocity prior. Despite this difference as well as slight differences in the observational scenarios, the confidence intervals on subhalo mass estimation are very similar to our results for LSST-era data.
Erkal & Belokurov (2015b) also estimated the minimum detectable subhalo mass based on individual observables, as a function of the error of that observable $\sigma_{\overline{X}}$. They similarly find that the shape of the stream on the sky (or $\overline{\Delta\theta}$) is generally the most discerning observable, with the potential to probe subhalos down to $\sim 10^{6}\mathrm{M}_{\odot}$. Our work provides a detailed statistical study to support this statement, while also accounting for different stream distances, densities, and dispersion.
@@
Another important work estimating subhalo detection in stellar streams is Drlica-Wagner et al. (2019). Focusing on detectability of a gap, they obtain the minimum detectable subhalo mass as a function of the stream surface brightness, which implicitly decides the stream density. While this work considered different stream distances and surface brightness, it did not consider stream width, which is another important factor in detectability. For a thin stream, the results are similar. Considering a stream with $\mu=31.5$ mag/arcsec2, $\sigma_{z}=20$ pc and $r_{0}=20$ kpc, the corresponding parameters are $\sigma_{\theta}=0.06^{\circ}$ and $\lambda=222$ deg-1. Plugging this into our Eq. 27, we find $M_{\mathrm{sh}}^{\mathrm{min}}=3.3\times 10^{5}~{}\mathrm{M}_{\odot}$ at 95% CL for LSST-10 year. This is somewhat smaller than their result of $10^{6}~{}\mathrm{M}_{\odot}$. However, that work requires a gap present at $5\sigma$ which is a higher threshold, and does not use other observables.
All of the above have focused on detectability of a single strong impact. Previous studies have estimated the number of strong impacts for a given stream (Erkal et al., 2016; Barry et al., 2023; Menker & Benson, 2024; Adams et al., 2024), which has typically led to $O(1)$ detectable gaps. For instance, recent works (Menker & Benson, 2024; Adams et al., 2024) have evaluated the distribution of gap sizes using semi-analytic models of structure formation. Menker & Benson (2024) uses the same analytic model of gap evolution for circular orbits, but improves upon it by numerically calculating velocity kicks for more realistic DM subhalo potentials. Meanwhile, Adams et al. (2024) goes beyond circular orbits and calculates velocity kicks using a 1D model of the stream, and then simulates stronger stream-subhalo encounters using gala. An interesting future direction is to evaluate the number of detectable impacts accounting for the properties of the stream, observables beyond density perturbations, as well as the DM subhalo population in a cosmological setting.
@@ 7 Conclusions
In this work, we have studied the detectability of DM subhalo impacts on stellar streams, taking into account stream properties such as width, distance, and stream density, as well as different observational scenarios. We have performed our study under the assumption of circular orbits for the stream, using the analytic model from Erkal & Belokurov (2015b) to enable fast simulation of impacts. Our statistical framework allows us to estimate the minimum detectable subhalo mass and the confidence interval for the detected subhalo mass in a given stream. One of our main results is a fitting formula, Eq. 27, that estimates the minimum detectable subhalo mass as a function of stream properties and for various observational scenarios. We applied this formula to 50 confirmed streams in the Milky Way with reported stellar mass, sufficient angular length, and sufficient stellar density. This led to a ranked list which identifies promising candidates in terms of DM subhalo impact detectability, as shown in Tab. 5. The following are our main conclusions:
@@
  • Having applied Eq. 27 to known streams in the Milky Way, we have identified promising candidate streams like C-12, ATLAS-Aliqa Uma, 300S, NGC 6397, and GD-1 which have the best sensitivity to low mass subhalos (see Tab. 5 and Sec. 5). While some of the most promising streams have been studied in detail for DM impacts already (Pal-5, GD-1 in particular), there are many others which have received more limited attention. This work adds motivation for more detailed follow up analyses and observations.
Though our results are based on an analytic model, which has a number of limitations, they compare very well to more sophisticated estimates done on some real streams, as discussed in Sec. 6.2. A thorough analysis of individual streams is much more expensive, but can take into account the actual stream geometry, the Milky Way potential including baryonic perturbers, the actual position of the observer, and other effects including background sources of perturbations and sources of errors. We briefly discussed potential generalizations in Sec. 6.1 and the method of orbit integration in App. B, which can take into account real stream orbits and a more general Milky Way potential, but leave the other details to future work. Ideally, one would eventually like to probe the DM subhalo distribution in the Milky Way by testing the expected impact signals against observations. Sensitivity to the distribution will be improved by increasing the effective exposure, and multiple streams can be considered in the analysis together for this purpose. Our work in this paper provides a first step to quickly assess the discovery threshold of subhalo impacts for a general stream in the Milky Way.
@@ Acknowledgements
Software: Python (Van Rossum & Drake, 2009), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), astropy (Robitaille et al., 2013; Price-Whelan et al., 2018; Collaboration et al., 2022), jupyter (Kluyver et al., 2016), matplotlib (Hunter, 2007), gala (Price-Whelan, 2017).
@@ Appendix A Distribution of test statistic
Figure 25: Top: $q_{0}$ distribution for different stream properties and observational scenarios. The first row is for the default GD-1 like stream, while in subsequent rows we vary the thickness (second row), distance (third row), and density (fourth row). The columns correspond to two observational scenaiors (DESI + Gaia and Via + Gaia + LSST10) at mass points $M_{\mathrm{sh}}^{\mathrm{true}}=0$ (no subhalo impact) and $M_{\mathrm{sh}}^{\mathrm{true}}=10^{8}~{}\mathrm{M}_{\odot}$ (high subhalo mass). Bottom: The corresponding distribution for $\hat{M}_{\mathrm{sh}}$, the subhalo mass that maximizes the likelihood function, with the same set of settings as top.

### v3

@@ 5 Application to known streams
Figure 19: Sky plots for individual streams ranked by detectability (left to right, top to bottom) under Via + Gaia + LSST10.
Figure 20: Visualized stream catalog, where color scale indicates the minimum detectable subhalo mass, $M_{\mathrm{sh}}^{\mathrm{min}}$, with respect to the best observational scenario. The stream properties of stellar density $\lambda$, width $\sigma_{\theta}$, and distance to observer $r_{0}$ all play a role in determining $M_{\mathrm{sh}}^{\mathrm{min}}$. The six streams with lowest $M_{\mathrm{sh}}^{\mathrm{min}}$ are labeled.
Figure 21: Improvement from DESI + Gaia to Via + Gaia + LSST in $\log_{10}M_{\mathrm{sh}}^{\mathrm{min}}/\mathrm{M}_{\odot}$ as functions of stream width (left), stream distance (middle) and stream density (right). The color scale indicates the minimum detectable subhalo mass, $M_{\mathrm{sh}}^{\mathrm{min}}$, with respect to the best observational scenario.
@@ 6.1 Generalization to realistic streams
Circular orbits and observer location – We expect that some of the largest differences between our analysis and more realistic models will arise in cases of large variation of heliocentric and/or Galactocentric distance along the stream orbit. In our study, streams are moving on circular orbits and observers are located at the Galactic Center. Accounting for eccentric orbits and placing the observer at the Sun will change the shape of subhalo-induced perturbations, their time dependence, as well as the relative importance of different observables. For a stream with a large variation in heliocentric distance along its length, a subhalo impact will be compressed to a shorter angular arc from the point of view of the observer, leading to a diminished detectability. Subhalo impacts will also be compressed and stretched depending on whether impacts occurred closer to pericenter or apocenter, as well as whether the stream today is observed near pericenter or apocenter. This will lead to much stronger dependence of subhalo detectability on impact time. We expect streams with relatively constant heliocentric and Galactocentric distances along the stream to have the most similar detectability as our idealized study.
Another effect in going from circular orbit to eccentric orbit is the increased fluctuations across angular bins caused by periodic stripping and epicyclic motion (Küpper et al., 2012). This introduces additional intrinsic fluctuations across bins, which breaks our assumption of a constant dispersion along the stream. We will leave the details of how to incorporate this effect into the statistical test for future work.
Furthermore, throughout the paper, we have been assuming that we know the progenitor orbit perfectly with zero uncertainty. However, the orbit depends on the Milky Way potential, and in practice stream tracks are used to constrain the potential (Koposov et al., 2010; Bonaca et al., 2014; Sanders, 2014). This introduces additional uncertainties which can reduce the significance of a signal, although this could be mitigated by simultaneous fits to the orbits of multiple streams (Reino et al., 2021). Another issue arises when there has been a massive impact which causes a perturbation extending over the entire stream, which can be challenging to distinguish from an alternative stream orbit. This mostly affects the detectability in more massive subhalo impacts ($\gtrsim 10^{8}~\mathrm{M}_{\odot}$ depending on the stream length), but has less of an effect on our results on the minimum detectable subhalo mass.
@@
Milky Way potential – The actual Milky Way potential is much more complicated than a simple logarithmic potential used here. Baryonic structures like the bar (Pearson et al., 2017), spiral arms, as well as giant molecular clouds (GMC) (Amorisco et al., 2016) and globular clusters (GC) (Erkal et al., 2017; Ferrone et al., 2025), could potentially cause perturbations in the streams. See also Banik and Bovy (2019), which considered the effect of all these structures on Pal-5. This will further raise the threshold of detectability, beyond the ideal case shown here. At the same time, kinematic information like radial velocity and proper motions could help distinguish between the impact of baryonic structures and DM subhalos (Bonaca et al., 2020; Hilmi et al., 2024; Price-Whelan et al., 2016).
@@
Stream model – Our stream model uses position/velocity dispersions assuming a particle spray model, where the parameters governing the dispersions are taken from Fardal et al. (2015). We investigated another particle spray model, based on a newer suite of simulations in Chen et al. (2024). This model includes initial radial velocity dispersions for the released particles and generates a stream with twice the $\sigma_{v_{r}}$ and $\sigma_{v_{t}}$ compared to a stream generated using Fardal et al. (2015), holding the angular width $\sigma_{\theta}$ fixed. For our default geometry where the subhalo moves along the $\hat{z}$ direction, this difference does not change the estimated detectability in a significant way, because the shift in the $z$ direction, $\overline{\Delta\theta}$, is the most effective observable in subhalo detection (see Fig. 9). The dispersion for this observable is governed primarily by the directly observed quantity $\sigma_{\theta}$. However, for more generic subhalo impact directions like $\hat{w}\sim\hat{z}+\hat{r}$ as in the third panel of Fig. 16, $\overline{\Delta v_{r}}$ also plays an important role. In this case, the minimum detectable subhalo mass will be higher by $\sim 0.2$ dex when adopting the particle spray model from Chen et al. (2024), due to its higher $\sigma_{v_{r}}$.
In addition to affecting the dispersions, the dynamics of progenitor disruption could induce perturbations in the streams and further raise the threshold for detectability. However, these effects are expected to mainly be in the region near the progenitor.
@@
Rather than fixing $r_{s}(M_{\mathrm{sh}})$, the degeneracy in mass and velocity can also be broken by using other subhalo potentials (such as Hernquist or NFW) and/or using the method of orbit integration, where we model the perturbed stream using the perturbed orbit of the progenitor (more details are discussed in App. B). The method of orbit integration does not rely on the impulse approximation, thus effectively breaking the degeneracy of the analytic model. We then evaluated how expected confidence intervals change when the $r_{s}(M_{\mathrm{sh}})$ relationship is removed and the method of orbit integration and/or Hernquist potential is used. For the Hernquist potential, the CDM mass-radius relation follows (Erkal et al., 2016):
@@
Fig. 22 shows a comparison of the modified approaches with our default approach (Analytic + Plummer + $r_{s}(M_{\mathrm{sh}})$). In all cases, we evaluated the test statistic on the Asimov data set. We find that using orbit integration generally results in tighter confidence intervals, even when $r_{s}$ is a free parameter of the fit. This indicates that our assumption of $r_{s}(M_{\mathrm{sh}})$ is not overly restrictive, and in fact the model likely underestimates the detectability compared to orbit integration. The results also show that Plummer and Hernquist potentials give similar signal strengths, as long as consistent $r_{s}$ values are used (see also the lower right panel of Fig. 27). Assessing the impact on the minimum detectable subhalo mass is much more expensive computationally, since it requires many more statistical realizations, and therefore we reserve this type of analysis for dedicated studies of individual streams.
Figure 22: Comparison of the test statistic $t(M_{\mathrm{sh}})$, as defined in Eq. 28, using the analytic model and using an orbit integration method. The dataset is generated from $M_{\mathrm{sh}}^{\mathrm{true}}=10^{8}~\mathrm{M}_{\odot}$. The stream has properties similar to GD-1 ($\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~\mathrm{kpc}$, $\lambda=100~\mathrm{deg}^{-1}$) and the observational scenario is LSST 10 year sensitivity.
More generally, the orbit integration method can be used to account for many of the concerns above. This would allow us to go beyond circular orbits and adopt a more general Milky Way potential, including known structures like the bar, spiral arms, and the Large Magellanic Cloud. It also provides an intermediate approach to modeling perturbed streams, which avoids orbit integration of the entire stream. It is a good approximation for CDM subhalos and again Gaussian noise can be added to account for dispersion. However, given that this approach is still much slower than the analytic model, we reserve this for future work.
@@ 6.2 Comparisons with previous work
Our work draws on Erkal and Belokurov (2015b), which derived the analytic model. That work focused on parameter estimation from a subhalo impact rather than the minimum detectable subhalo. Using a Bayesian framework, the 6 observables from the analytic model are used to infer all the parameters of the subhalo impact at the same time. To address the subhalo velocity-mass degeneracy of the analytic model discussed above, they imposed a subhalo velocity prior rather than a mass-radius relationship. The confidence bands are then highly dependent on the velocity prior. Despite this difference as well as slight differences in the observational scenarios, the confidence intervals on subhalo mass estimation are very similar to our results for LSST-era data.
Erkal and Belokurov (2015b) also estimated the minimum detectable subhalo mass based on individual observables, as a function of the error of that observable $\sigma_{\overline{X}}$. They similarly find that the shape of the stream on the sky (or $\overline{\Delta\theta}$) is generally the most discerning observable, with the potential to probe subhalos down to $\sim 10^{6}\mathrm{M}_{\odot}$. Our work provides a more robust result by performing a statistical test considering all 6 observables at the same time, while also accounting for different stream distances, densities, and dispersion.
@@
Another important work estimating subhalo detection in stellar streams is Drlica-Wagner et al. (2019). Focusing on detectability of a gap, they obtain the minimum detectable subhalo mass as a function of the stream surface brightness, which implicitly decides the stream density. While this work considered different stream distances and surface brightness, it did not consider stream width, which is another important factor in detectability. For a thin stream, the results are similar. Considering a stream with $\mu=31.5$ mag/arcsec2, $\sigma_{z}=20$ pc and $r_{0}=20$ kpc, the corresponding parameters are $\sigma_{\theta}=0.06^{\circ}$ and $\lambda=222$ deg-1. Plugging this into our Eq. 27, we find $M_{\mathrm{sh}}^{\mathrm{min}}=3.3\times 10^{5}~\mathrm{M}_{\odot}$ at 95% CL for LSST 10 year. This is somewhat smaller than their result of $10^{6}~\mathrm{M}_{\odot}$. However, that work requires a gap present at $5\sigma$ which is a higher threshold, and does not use other observables.
All of the above have focused on detectability of a single strong impact. Previous studies have estimated the number of strong impacts for a given stream (Erkal et al., 2016; Barry et al., 2023; Menker and Benson, 2024; Adams et al., 2024), which has typically led to $O(1)$ detectable gaps. For instance, recent works (Menker and Benson, 2024; Adams et al., 2024) have evaluated the distribution of gap sizes using semi-analytic models of structure formation. Menker and Benson (2024) uses the same analytic model of gap evolution for circular orbits, but improves upon it by numerically calculating velocity kicks for more realistic DM subhalo potentials. Meanwhile, Adams et al. (2024) goes beyond circular orbits and calculates velocity kicks using a 1D model of the stream, and then simulates stronger stream-subhalo encounters using gala. An interesting future direction is to evaluate the number of detectable impacts accounting for the properties of the stream, observables beyond density perturbations, as well as the DM subhalo population in a cosmological setting.
@@ 7 Conclusions
In this work, we have studied the detectability of DM subhalo impacts on stellar streams, taking into account stream properties such as width, distance, and stream density, as well as different observational scenarios. We have performed our study under the assumption of circular orbits for the stream, using the analytic model from Erkal and Belokurov (2015b) to enable fast simulation of impacts. Our statistical framework allows us to estimate the minimum detectable subhalo mass and the confidence interval for the detected subhalo mass in a given stream. One of our main results is a fitting formula, Eq. 27, that estimates the minimum detectable subhalo mass as a function of stream properties and for various observational scenarios. We applied this formula to 50 confirmed streams in the Milky Way with reported stellar mass, sufficient angular length, and sufficient stellar density. This led to a ranked list which identifies promising candidates in terms of DM subhalo impact detectability, as shown in Tab. 5. The following are our main conclusions:
@@
  • Having applied Eq. 27 to known streams in the Milky Way, we have identified promising candidate streams like C-12, ATLAS-Aliqa Uma, 300S, NGC 6397, Palomar 5 and GD-1 which have the best sensitivity to low mass subhalos (see Tab. 5 and Sec. 5). While some of the most promising streams have been studied in detail for DM impacts already (Pal-5, GD-1 in particular), there are many others which have received more limited attention. This work adds motivation for more detailed follow up analyses and observations.
Though our results are based on an analytic model, which has a number of limitations, they compare very well to more sophisticated estimates done on some real streams, as discussed in Sec. 6.2. A thorough analysis of individual streams is much more expensive, but can take into account the actual stream geometry, the Milky Way potential including baryonic perturbers, the actual position of the observer, and other effects including background sources of perturbations and sources of errors. We briefly discussed potential generalizations in Sec. 6.1 and the method of orbit integration in App. B, which can take into account real stream orbits and a more general Milky Way potential, but leave the other details to future work.
Furthermore, we emphasize that the stream ranking on the minimum detectable subhalo mass we calculated in this work is not the complete picture for the most promising streams. The detection prospects also depend on the expected rate of subhalo impacts, which potentially depends on stream length and stream age, as well as baryonic backgrounds. By incorporating all these effects, one would eventually probe the DM subhalo distribution by testing the expected number of detectable impacts against observations. Sensitivity to the distribution will be improved by increasing the effective exposure, and multiple streams can be considered in the analysis together for this purpose. Our work in this paper provides a first step to quickly assess the discovery threshold of subhalo impacts for a general stream in the Milky Way, and the remained steps will be done in future work.
@@ Acknowledgements
Software: Python (Van Rossum and Drake, 2009), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), astropy (Robitaille et al., 2013; Price-Whelan et al., 2018; Collaboration et al., 2022), jupyter (Kluyver et al., 2016), matplotlib (Hunter, 2007), gala (Price-Whelan, 2017).
@@ Appendix A Distribution of test statistic
Figure 26: Top: $q_{0}$ distribution for different stream properties and observational scenarios. The first row is for the default GD-1 like stream, while in subsequent rows we vary the thickness (second row), distance (third row), and density (fourth row). The columns correspond to two observational scenarios (DESI + Gaia and Via + Gaia + LSST10) at mass points $M_{\mathrm{sh}}^{\mathrm{true}}=0$ (no subhalo impact) and $M_{\mathrm{sh}}^{\mathrm{true}}=10^{8}~\mathrm{M}_{\odot}$ (high subhalo mass). Bottom: The corresponding distribution for $\hat{M}_{\mathrm{sh}}$, the subhalo mass that maximizes the likelihood function, with the same set of settings as top.